OpenStudy (bailey):

Use the quadratic formula to solve the equation..4w2+20w+25=0

OpenStudy (anonymous):

divide by 2

OpenStudy (anonymous):

wait2..its wrong

OpenStudy (bailey):

Ahh nevermind.

OpenStudy (bailey):

I got it

OpenStudy (anonymous):

(2x + 5)(2x + 5) = 0 x = -5/2

OpenStudy (anonymous):

w=-b+\[\sqrt{b ^{2}-4ac}\]2a

OpenStudy (bailey):

x2-5x+14=0

OpenStudy (anonymous):

w=[-20+-surd(20^2 -4(4)(25))]/2(4)

OpenStudy (anonymous):

i dun no how to use this equation symbol

OpenStudy (bailey):

Blexting! help

OpenStudy (anonymous):

blexting: that was factorization method

OpenStudy (bailey):

Thats right though

OpenStudy (bailey):

3x2 -4 +11x solve equation

OpenStudy (anonymous):

ya its right but you have to answer the qstn by using quadratic formula.thas what the qstn wanted you to do

OpenStudy (bailey):

Thats true i dont care, we dont have to use that formula. I just want the right answer

OpenStudy (anonymous):

(3x-1)(x+4)=0 x=1/3,x=-4

OpenStudy (anonymous):

Sorry.. I was away.. do you need anything else?

OpenStudy (anonymous):

x^2 - 5x + 14 = 0 quadratic formula 5 +- sqrt (25 - 4 (1) (14) --------------------- 2 5 +- sqrt(25 - 56) --------------- 2 5 +- sqrt(-31) ------------- 2 5 +- i sqrt(31) -------------- 2